# An algebraic model for rational excisive functors

### v1

@@ 6. Splitting rational excisive functors
$\simeq{\mathsf{Mod}}_{\mathbb{Q}}(D_{d}h_{\mathbb{S}})\times\mathsf{Exc}_{d-1}({{\mathsf{Sp}}},{{\mathsf{Sp}}})_{\mathbb{Q}},$
since $D_{d}h_{\mathbb{S}}$ inherits the structure of a commutative algebra object in ${\mathsf{Fun}}({{\mathsf{Sp}}},{{\mathsf{Sp}}})_{\mathbb{Q}}$ through the equivalence $D_{d}h_{\mathbb{S}}\simeq P_{d}h_{\mathbb{S}}[e_{d}^{-1}]$.
Through the equivalent descriptions of the splitting of $\mathsf{Exc}_{d}({{\mathsf{Sp}}},{{\mathsf{Sp}}})_{\mathbb{Q}}$ the projection onto the second factor corresponds to the functor
$P_{d-1}:\mathsf{Exc}_{d}({{\mathsf{Sp}}},{{\mathsf{Sp}}})_{\mathbb{Q}}\longrightarrow\mathsf{Exc}_{d-1}({{\mathsf{Sp}}},{{\mathsf{Sp}}})_{\mathbb{Q}},$
@@
$\mathsf{Exc}_{d}({{\mathsf{Sp}}},{{\mathsf{Sp}}})_{\mathbb{Q}}\simeq\mathsf{Homog}_{d}({{\mathsf{Sp}}},{{\mathsf{Sp}}})_{\mathbb{Q}}\times\mathsf{Exc}_{d-1}({{\mathsf{Sp}}},{{\mathsf{Sp}}})_{\mathbb{Q}}.$
The splitting as a product of $\infty$-categories of homogenous functors follows by reverse induction and the observation that there is an equivalence $\mathsf{Homog}_{1}({{\mathsf{Sp}}},{{\mathsf{Sp}}})_{\mathbb{Q}}\simeq\mathsf{Exc}_{1}({{\mathsf{Sp}}},{{\mathsf{Sp}}})_{\mathbb{Q}}$ since all functors are assumed reduced. ∎
The proof of 6.4 has the following immediate corollary, which amounts to the observation that rational homogeneous functor $D_{d}h_{\mathbb{S}}$ admits the structure of a commutative algebra object in ${\mathsf{Fun}}({{\mathsf{Sp}}},{{\mathsf{Sp}}})_{\mathbb{Q}}$. There is a priori no reason for this result to hold integrally, since $D_{d}$ need not be a symmetric monoidal functor.
@@
$\mathsf{Homog}_{d}({{\mathsf{Sp}}},{{\mathsf{Sp}}})_{\mathbb{Q}}\simeq{\mathsf{Mod}}_{\mathbb{Q}}(D_{d}h_{\mathbb{S}}),$
@@
It is also possible to prove the identification of the $(1-e_{d})$-local piece of  6.4 by considering the adjunction
[…]
and applying the Left Localization Principle of Pol and Williamson [PW20, Theorem 3.13] with respect to $(1-e_{d})$-localization.
@@ 7. An algebraic model for rational excisive functors
The main theorem of Goodwillie calculus is the classification of homogeneous functors. To every functor, $F:{{\mathsf{Sp}}}\to{{\mathsf{Sp}}}$ Goodwillie assigns a symmetric sequence $\partial_{\ast}F=\{\partial_{d}F\}_{d\in\mathbb{N}}$ of spectra, in which we call the $d$-th term $\partial_{d}F$, the $d$-th derivative of $F$. When the input functor is $d$-homogeneous, Goodwillie shows that it is completely determined by its $d$-th derivative. By [Goo03, Theorem 3.5 and §5], there is an equivalence
$\mathsf{Homog}_{d}({{\mathsf{Sp}}},{{\mathsf{Sp}}})\simeq{{\mathsf{Sp}}}^{B\Sigma_{d}}$
between the $\infty$-category of $d$-homogeneous endofunctors of spectra and the $\infty$-category of spectra with an action of $\Sigma_{d}$, the symmetric group on $d$ letters. Applying this classification of homogeneous functors to the splitting of 6.4 we obtain the following.
@@
$\mathsf{Exc}_{d}({{\mathsf{Sp}}},{{\mathsf{Sp}}})_{\mathbb{Q}}\simeq\prod_{1\leq i\leq d}{{\mathsf{Sp}}}_{\mathbb{Q}}^{B\Sigma_{i}}$
@@
Shipley’s [Shi07] equivalence between $HR$-module spectra and chain complexes of $R$-modules together with the identification of $H\mathbb{Q}$-module spectra with rational spectra induce the algebraic model of B upon the observation that there is an induced equivalence
@@
$\mathsf{Exc}_{d}({{\mathsf{Sp}}},{{\mathsf{Sp}}})_{\mathbb{Q}}\simeq\prod_{1\leq i\leq d}{\mathsf{Ch}}(\mathbb{Q}[\Sigma_{i}])$
of $\infty$-categories. We call the right-hand side an algebraic model of the $\infty$-category of rational $d$-excisive functors from spectra to spectra.

### v2

@@ 6. Splitting rational excisive functors
$\simeq{\mathsf{Mod}}_{\mathbb{Q}}(D_{d}h_{\mathbb{S}})\times\mathsf{Exc}_{\ast}^{d-1}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})_{\mathbb{Q}},$
since $D_{d}h_{\mathbb{S}}$ inherits the structure of a commutative algebra object in ${\mathsf{Fun}}_{\ast}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})_{\mathbb{Q}}$ through the equivalence $D_{d}h_{\mathbb{S}}\simeq P_{d}h_{\mathbb{S}}[e_{d}^{-1}]$.
Through the equivalent descriptions of the splitting of $\mathsf{Exc}_{\ast}^{d}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})_{\mathbb{Q}}$ the projection onto the second factor corresponds to the functor
$P_{d-1}:\mathsf{Exc}_{\ast}^{d}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})_{\mathbb{Q}}\longrightarrow\mathsf{Exc}_{\ast}^{d-1}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})_{\mathbb{Q}},$
@@
$\mathsf{Exc}_{\ast}^{d}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})_{\mathbb{Q}}\simeq\mathsf{Homog}^{d}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})_{\mathbb{Q}}\times\mathsf{Exc}_{\ast}^{d-1}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})_{\mathbb{Q}}.$
The splitting as a product of $\infty$-categories of homogenous functors follows by reverse induction and the observation that there is an equivalence $\mathsf{Homog}^{1}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})_{\mathbb{Q}}\simeq\mathsf{Exc}_{\ast}^{1}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})_{\mathbb{Q}}$ since all functors are assumed reduced. ∎
The proof of ˜6.4 has the following immediate corollary, which amounts to the observation that rational homogeneous functor $D_{d}h_{\mathbb{S}}$ admits the structure of a commutative algebra object in ${\mathsf{Fun}}_{\ast}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})_{\mathbb{Q}}$.
@@
$\mathsf{Homog}^{d}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})_{\mathbb{Q}}\simeq{\mathsf{Mod}}_{{\mathsf{Fun}}_{\ast}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})_{\mathbb{Q}}}(D_{d}h_{\mathbb{S}}),$
@@
There is a priori no reason for this result to hold integrally: $D_{d}$ is not a symmetric monoidal endofunctor of ${\mathsf{Fun}}_{\ast}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})$ since $D_{d}$ does not preserve the symmetric monoidal unit of ${\mathsf{Fun}}_{\ast}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})$. The functor $D_{d}h_{\mathbb{S}}$ is an idempotent coalgebra object in ${\mathsf{Fun}}_{\ast}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})$ since the colocalization $D_{d}:\mathsf{Exc}_{\ast}^{d}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})\to\mathsf{Homog}^{d}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})$ is smashing, i.e., given by Day convolution with $D_{d}h_{\mathbb{S}}$. In the rational setting, the identification of $D_{d}h_{\mathbb{S}}$ with the localization of $P_{d}h_{\mathbb{S}}$ at the idempotent $e_{d}$ induces an idempotent algebra structure on $D_{d}h_{\mathbb{S}}$ which is compatible with the idempotent coalgebra structure. This algebra structure cannot lift to an integral algebra structure as if it did, the splitting of ˜6.4 would also hold integrally.
@@ 7. An algebraic model for rational excisive functors
One of the major theorems of Goodwillie calculus is the classification of homogeneous functors. To every functor, $F:{{\mathsf{Sp}}}\to{{\mathsf{Sp}}}$ Goodwillie assigns a symmetric sequence $\partial_{\ast}F=\{\partial_{d}F\}_{d\in\mathbb{N}}$ of spectra, in which we call the $d$-th term $\partial_{d}F$, the $d$-th derivative of $F$. When the input functor is $d$-homogeneous, Goodwillie shows that it is completely determined by its $d$-th derivative. By [Goo03, Theorem 3.5 and §5], there is an equivalence
$\mathsf{Homog}^{d}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})\simeq{{\mathsf{Sp}}}^{B\Sigma_{d}}$
between the $\infty$-category of $d$-homogeneous endofunctors of spectra and the $\infty$-category of spectra with an action of $\Sigma_{d}$, the symmetric group on $d$ letters. Applying this classification of homogeneous functors to the splitting of ˜6.4 we obtain the following.
@@
$\mathsf{Exc}_{\ast}^{d}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})_{\mathbb{Q}}\simeq\prod_{1\leq i\leq d}{{\mathsf{Sp}}}_{\mathbb{Q}}^{B\Sigma_{i}}$
@@
Shipley’s [Shi07] equivalence between $HR$-module spectra and chain complexes of $R$-modules together with the identification of $H\mathbb{Q}$-module spectra with rational spectra induce the algebraic model of ˜B upon the observation that there is an induced equivalence
@@
$\mathsf{Exc}_{\ast}^{d}({{\mathsf{Sp}}}^{\omega},{{\mathsf{Sp}}})_{\mathbb{Q}}\simeq\prod_{1\leq i\leq d}{\mathsf{Ch}}(\mathbb{Q}[\Sigma_{i}])$
of $\infty$-categories.
We call the right-hand side an algebraic model of the $\infty$-category of rational $d$-excisive functors from spectra to spectra.